OpenStudy (anonymous):

pls help me solve by using substitution method: 5x+35y= -10 15x +5y = 170 I am so tired I cant think

OpenStudy (anonymous):

(5x+35y=-10)-3 15x+5y=170 -15x-105y=30 15x+5y=170 -100y=200 y=-2 solve for x

OpenStudy (anonymous):

how did you get that ? the -15x-105y=30?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

oh you said substitution, i used elimination, sorry

OpenStudy (anonymous):

lol thats ok :)

hero (hero):

5x + 35y = -10 ----> x + 7y = -2 ---> y = (-2 - x)/7 15x + 5y = 170 ----> 3x + y = 34 ---> y = 34 - 3x y = y 7(34 - 3x) = -2 - x 238 - 21x = -2 - x 240 = 20x 240/20 = x 12 = x -2 = y

hero (hero):

Good luck trying to figure that out. :D

OpenStudy (anonymous):

thx :)preciate it